# An improved description of nuclear charge radii: Global trends beyond $N=28$ shell closure

### v1

@@ IV Conclusions
The neutron-proton ($np$) correlation originating from the unpaired neutron and proton are taken into account in describing nuclear charge radii. The systematic study of charge radii for nuclei with proton number $Z=19$-$29$ isotopes are performed. The odd-even staggering (OES) and the inverted parabolic-like behaviors in charge radii are reproduced well along potassium and calcium isotopes. Although the NL3, PK1 and $*{}^{*}$ effective forces give the similar trend, the PK1 parameter set systematically underestimates the results. Beyond $N=28$ shell closure, the sharp increases are reproduced remarkably well. Especially the charge radii of nickel and copper isotopes can be reproduced against the latest experimental data well. It can be found that the increasing trend of charge radii beyond $N=28$ shell closure is almost independent of proton number, which is in accord with Ref. Kortelainen et al. (2022). Our calculations can give the shrunk phenomenon of charge radii of Sc isotopes at $N=20$, but the trend is underestimated.
@@
In general, sudden drop in nucleon separation energies, a larger $2^{+}_{1}$ excitation energy and a smaller transition probability as compared to the neighbor isotones (or isotopes) are commonly considered as manifestations of the magic character of a nucleus. Along isotopic chains, unexpected increases of charge radii are not observed across neutron number $N=32$ Koszorús et al. (2019, 2021). The shell effect leads to the kink in the isotopic chains, where charge radii vary smoothly closing to neutron magic numbers, and then the sudden increase occurs after crossing filled shells. In our calculations, the calculated results monotonically increase with the increasing neutron number along $Z=19$-$29$ isotopic chains. Therefore, it is difficult to identify the new magicity of nuclei with neutron number $N=32$ and $34$ along a long isotopic chain.
@@
However, a high interest is that the local variations of charge radii along isotonic chains. In our calculations, the abrupt change of nuclear charge radii across $Z=28$ is evidently shown along $N=28$ isotonic chain, but this could be weakened across $Z=20$. Along $N=32$ and $34$ isotones, the abrupt increases of charge radii are also predicted across the proton number $Z=22$, but later with a lower slope. In potassium and calcium isotopes, $N=32$ possesses the magicity in the view of relative stable binding properties. However, the shrinking phenomena cannot be observed in charge radii of nuclei with featuring $N=32$. Our results suggest that the trend of changes in charge radii is almost linear from potassium $(Z=19)$ to titanium $(Z=22)$. This is consistent with the Ref. Reiter et al. (2018) where the signature of $N=32$ sub-shell closure is not identified across V ($Z=23$) and even higher proton numbers. The same scenario can also be observed along $N=34$ isotonic chain, but more available results are needed. An indication of the robustness of a new shell closure may be identified from these aspects.
@@
An available description of the nuclear size can provide access to understand new physics beyond standard model Safronova et al. (2018) and serve as input quantities in astrophysical study Arnould and Goriely (2020). As demonstrated in Ref. Perera and Afanasjev (2023), proton-neutron interaction is responsible for the fine structure of charge radii. In our calculations, the unpaired neutron-proton correlation is tentatively incorporated into the root-mean square charge radii formula. The overestimated odd-even staggering in charge radii are definitely improved by this modified formula. It is known that the linear relationship between the difference of the charge radii of mirror nuclei and nuclear symmetry energy can be used to inspect the isovector forces Novario et al. (2023); Yang and Piekarewicz (2018). Therefore, more available models are needed in describing nuclear size and more data about the charge radii of mirror partner nuclei are required in experiments.
Acknowledgements
@@
This work is supported in part by the National Natural Science Foundation of China under Grants No. 12047513, No. 12135004, No. 11635003, No. 11961141004, No. 11025524, No. 11161130520, the National Basic Research Program of China under Grant No. 2010CB832903. X. J. is grateful for the support of the National Natural Science Foundation of China under Grants No. 11705118. L.-G. C. is grateful for the support of the National Natural Science Foundation of China under Grants No. 12275025, No. 11975096 and the Fundamental Research Funds for the Central Universities (2020NTST06).

### v2

@@ IV Conclusions
The neutron-proton ($np$) correlation originating from the unpaired neutron and proton are taken into account in describing nuclear charge radii. The systematic study of charge radii for nuclei with proton numbers $Z=19$-$29$ isotopes are performed. The odd-even staggering (OES) and the inverted parabolic-like behaviors in charge radii are reproduced well along potassium and calcium isotopes. Although the NL3, PK1, and NL3 effective forces give the similar trend, the PK1 parameter set systematically underestimates the results. Beyond $N=28$ shell closure, the sharp increases are reproduced remarkably well. Especially the charge radii of nickel and copper isotopes can be reproduced against the latest experimental data well. It can be found that the increasing trend of charge radii beyond $N=28$ shell closure is almost independent of proton number, which is in accord with Ref. Kortelainen et al. (2022). Our calculations can give the shrunk phenomenon of charge radii of Sc isotopes at $N=20$, but the trend is underestimated.
@@
In general, sudden drop in nucleon separation energies, a larger $2^{+}_{1}$ excitation energy, and a smaller transition probability as compared to the neighbor isotones (or isotopes) are commonly considered as manifestations of the magic character of a nucleus. Along isotopic chains, unexpected increases of charge radii are not observed across neutron number $N=32$ Koszorús et al. (2019, 2021). The shell effect leads to the kink in the isotopic chains, where charge radii vary smoothly closing to neutron magic numbers, and then the sudden increase occurs after crossing filled shells. In our calculations, the calculated results monotonically increase with the increasing neutron number along $Z=19$-$29$ isotopic chains. Therefore, it is difficult to identify the new magicity of nuclei with neutron number $N=32$ and $34$ along a long isotopic chain.
@@
However, a high interest is the local variations of charge radii along isotonic chains. In our calculations, the abrupt change of nuclear charge radii across $Z=28$ is evidently shown along $N=28$ isotonic chain, but this could be weakened across $Z=20$. Along $N=32$ and $34$ isotones, the abrupt increases of charge radii are also predicted across the proton number $Z=22$, but the latter with a lower slope. In potassium and calcium isotopes, $N=32$ possesses the magicity in the view of relative stable binding properties. However, the shrinking phenomena cannot be observed in charge radii of nuclei with featuring $N=32$. Our results suggest that the trend of changes in charge radii is almost linear from potassium $(Z=19)$ to titanium $(Z=22)$. This is consistent with the Ref. Reiter et al. (2018) where the signature of $N=32$ sub-shell closure is not identified across V ($Z=23$) and nucleus with higher proton numbers. The same scenario can also be observed along $N=34$ isotonic chain, but more available results are needed. An indication of the robustness of a new shell closure may be identified from these aspects.
@@
An available description of the nuclear size can provide access to understand new physics beyond standard model Safronova et al. (2018) and serve as input quantities in astrophysical study Arnould and Goriely (2020). As demonstrated in Ref. Perera and Afanasjev (2023), proton-neutron interaction is responsible for the fine structure of charge radii. In our calculations, the unpaired neutron-proton correlation is tentatively incorporated into the root-mean-square charge radii formula. The overestimated odd-even staggering in charge radii are definitely improved by this modified formula. Although this modified formula can reproduce the local variations of nuclear charge radii, the self-consistently microscopic form is urgently required. It is known that the linear relationship between the difference of the charge radii of mirror nuclei and nuclear symmetry energy can be used to inspect the isovector forces Novario et al. (2023); Yang and Piekarewicz (2018). Therefore, more available models are needed in describing nuclear size and more data about the charge radii of mirror partner nuclei are required in experiments.
Acknowledgments
@@
This work is supported in part by the National Key R&\&&D Program of China under Grant No. 2023YFA1606401 and the National Natural Science Foundation of China under Grants No. 12047513, No. 12135004, No. 11635003, No. 11961141004, No. 11025524, No. 11161130520, the National Basic Research Program of China under Grant No. 2010CB832903. X. J. is grateful for the support of the National Natural Science Foundation of China under Grants No. 11705118. L.-G. C. is grateful for the support of the National Natural Science Foundation of China under Grants No. 12275025, No. 11975096 and the Fundamental Research Funds for the Central Universities (2020NTST06).